Ask your own question, for FREE!
Mathematics 10 Online
OpenStudy (anonymous):

If you perform the experiment, "Toss Three Coins", what is the probability of getting three tails? 1/8 3/8 1/2

OpenStudy (anonymous):

well the probability of the first being tails is 1/2, the second is 1/2, and the third is 1/2 the probability of getting all three tails would then be: 1/2 x 1/2 x 1/2 = 1/8

OpenStudy (chrisasl):

It's 1/8.. 1/2 for the 1st, 2nd and 3rd>> 1/2 * 1/2 * 1/2 = 1/8

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!